A diver springs upward with an initial speed of 1.92 m/s from a 3.0-m board. (a) Find the velocity with which he strikes the water. [Hint: When the diver reaches the water, his displacement is y = -3.0 m (measured from the board), assuming that the downward direction is chosen as the negative direction.] (b) What is the highest point he reaches above the water?
Answer #1

a)

From

V2=Vo2+2ad

V2=1.922+2(-9.81)(-3)=62.5464 ,

V=-7.91 m/s

when it strikes the water ,it is directed downward

b)

From

V2=Vo2+2a(H-ho)

when it reaches highest point V=0

=>0=1.922+2(-9.81)(H-3)

H=3.188 m